OpenStudy (anonymous):

I need help with this. Find ask real number solutions in the interval [0,2pi) that satisfy each equation. Round approximate answers to the nearest tenth. 2cos^2(2x)-8sin^2(x)cos^2(x)=-1 I don't see any identity substitutions or a way this factors. Hopefully I'm overlooking something.

OpenStudy (tommynaut):

I'm not sure what level of trig is required here but are you familiar with the identity cos2x = cos^2(x) - sin^2(x)? Knowing that, you can do some dirty maths with 2cos^2(2x), which is really just 2 times cos(2x) times cos(2x).

OpenStudy (anonymous):

I'm familiar with it. I guess I hadn't noticed that.

OpenStudy (tommynaut):

Are you able to go on from there?

OpenStudy (anonymous):

Maybe, I'm going to try.

OpenStudy (anonymous):

I was unable to solve it.

OpenStudy (tommynaut):

What did you get up to? It's quite a hard question really, either that or I did it a super weird way, but after expanding (cos(2x))^2 and putting it back into the equation, I was able to simplify things to get cos^2(x) sin^2(x) = 1/4, and then I used the trig identity sin2x = 2sinxcosx to simplify things even further

OpenStudy (anonymous):

I think I went a different way. so, you had 2[cos^2(x)-sin^2(x)][cos^2(x)-sin^2(x)]-8sin^2(x)cos^(x)=-1 Did you factor out a [cos^2(x)-sin^2(x)]?

OpenStudy (tommynaut):

Not quite. I just multiplied them together, using the fact that (a-b)^2 = a^2 - 2ab + b^2

OpenStudy (tommynaut):

It looks like a disgusting way of doing it (and it might not be the fastest way, idk) but we would end up with cos^4(x) + sin^4(x) which is actually just equal to 1

OpenStudy (tommynaut):

I mean, cos^4(x) + sin^4(x) in addition to a -2cos^2(x)sin^2(x)

OpenStudy (anonymous):

so

OpenStudy (anonymous):

so 2cos^4(x)-2cos^2(x)sin^2(x)+sin^4(x)-8sin^2(x)cos^2(x)=-1 2cos^4(x)+sin^4(x)-2cos^2(x)sin^2(x)-8sin^2(x)cos^2(x)=-1 I still don't see it.

OpenStudy (tommynaut):

You made a bit of an error. The 2 outside the cos^2(2x) applies to its entire expansion. So it would actually be: 2(cos^4(x) - 2cos^2(x)sin^2(x)+ sin^4(x)) -8sin^2(x)cos^2(x)=-1 2cos^4(x) - 4cos^2(x)sin^2(x)+ 2sin^4(x) -8sin^2(x)cos^2(x)=-1 Now we know that sin^2(x) + cos^2(x) = 1. The thing is, sin^4(x) + cos^4(x) = 1 too (you can go about proving that to yourself later if you like).

OpenStudy (anonymous):

You applied that identity after or before multiplying the two back in??

OpenStudy (tommynaut):

Well, before really, but I was just showing you what it should look like if you multiplied through by 2. In my own working I had it as 2(cos^4(x) - 2cos^2(x)sin^2(x)+ sin^4(x)) -8sin^2(x)cos^2(x)=-1 2(1 - 2cos^2(x)sin^2(x)) -8sin^2(x)cos^2(x)=-1

OpenStudy (tommynaut):

Actually to be honest I'm beginning to doubt that sin^4(x) + cos^4(x) is really equal to 1... sorry that's my bad :(

OpenStudy (anonymous):

1 times 1 is 1. I'm willing to use it for now. I can email a solution to my teacher tomorrow so she can check it.

OpenStudy (tommynaut):

Sorry but it's not actually equal to 1, I was getting mixed up with something else. I don't really have time to finish helping with this question, but if I was you I would write down every trig identity you know and try and see how you can apply a combination of them in the question to simplify it down into a situation where you only have 1 type of trig function. Then, you might be able to use a substitution like u = cos^2(x) and factorise your new u equation. Sorry about that.

OpenStudy (anonymous):

Thanks for trying.